# Bright and dark breathers on an elliptic wave in the defocusing mKdV equation

### v1

@@ 3.2. Characteristic polynomial for traveling waves
where all functions depend on only one variable which stands for the traveling wave coordinate $x+ct$. Since (3.12) is the linear algebraic system, admissible values of $\mu$ are found from the characteristic equation
$\left|\begin{array}[]{ll}4i\zeta^{3}+2i\zeta\phi^{2}-ic\zeta-\mu&4\zeta^{2}\phi-2i\zeta\phi^{\prime}+2\phi^{3}-\phi^{\prime\prime}-c\phi\\ 4\zeta^{2}\phi+2i\zeta\phi^{\prime}+2\phi^{3}-\phi^{\prime\prime}-c\phi&-4i\zeta^{3}-2i\zeta\phi^{2}+ic\zeta-\mu\end{array}\right|=0.$ (3.13)
@@
due to the parameterization (1.7). It follows from (3.12) that the quotient $\rho=q/p$ for the eigenfunction $\psi=(p,q)^{T}$ satisfies
Substituting (2.13) and (3.6) into (3.14) and using (2.2), we obtain
@@
where we used the first-order quadrature $(\wp^{\prime})^{2}=4(\wp-e_{1})(\wp-e_{2})(\wp-e_{3})$ for Weierstrass’ elliptic function $\wp$. It remains to prove that $\mu=-2\wp^{\prime}(a)$.
@@ 3.3. Explicit expression for $\mu$
First, we note from (1.16) that $p_{1}(x_{0})=q_{1}(x_{0})$ at
@@
where we have used expressions (2.4), (2.5), (2.6), (2.7), and (3.6). By using [16, 8.177],
@@
we obtain $\mu=-2\wp^{\prime}(a)$. Since $a=(\gamma+iK^{\prime})/\nu$, we use (1.22) and (2.1) to rewrite the expression for $\mu$ in the form
$\mu=-2\wp^{\prime}(a)=\frac{4\nu^{3}\operatorname{cn}(\nu a)\operatorname{dn}(\nu a)}{\operatorname{sn}^{3}(\nu a)}=-4\nu^{3}k^{2}\operatorname{sn}(\gamma)\operatorname{cn}(\gamma)\operatorname{dn}(\gamma),$ (3.16)
which yields (1.18). The explicit expression (3.16) enables a unique parameterization of the square root singularity in the definition (3.14) of $\mu$. We give explicit expressions for $\mu$ in the gaps $(\zeta_{2},\zeta_{1})$ and $(0,\zeta_{3})$ to show that the different branches of the square root are chosen in between the two gaps.
@@
due to (2.17).
For $\zeta\in(0,\zeta_{3})$, we write $\gamma=\delta+iK^{\prime}$ with $\delta\in(2\alpha,K)$. The expression (3.16) can be rewritten as
@@
since the dispersion relation (2.17) can be rewritten as
@@ 4. Breathers on the elliptic wave background
We use a linear combination of the two eigenfunctions in Theorem 1 in order to construct breathers by using the Darboux transformation (1.25). This leads to the expression (1.26), where we intend to show for the half-gap $(0,\zeta_{3})$ and the gap $(\zeta_{2},\zeta_{1})$ separately that the choice (1.30) produces a bounded and real-valued solution of the mKdV equation (1.1). Sections 4.1 and 4.2 report the corresponding results in the two gaps, from which breather solutions in Figures 34 and 56 were constructed.
@@ 4.1. Breathers in the half gap $(0,\zeta_{3})$
Figure 7. ${\rm Re}(\rho)$ versus $x$ (left) and ${\rm Im}(\rho)$ versus $x$ (right), where $\rho=q/p$ is the quotient of the two components of eigenfunctions $\psi=(p,q)^{T}$ of the spectral problem (3.7) with the elliptic potential $\phi$ for $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$. The spectral parameter $\zeta$ is defined by (3.4) with $\delta=\frac{1}{2}(2\alpha+K)$. The green line shows $\rho$ for the first solution (4.1) and the magenta line shows $\rho$ for he second solution (4.2).
For the numerical check, we plot in Figure 7 the quotient $\rho=q/p$ (real part is shown on the left panel and imaginary part is on the right panel) for the first solution (4.1) (green line) and the second solution (4.2) (magenta line). The quotient $\rho$ is computed in two different ways: by using the elliptic functions (3.15) with $\mu$ given by (3.18) and by using the explicit solutions (4.1) and (4.2). The difference between the two computational formulas is found within the machine precision error.
The breather solutions are obtained from (1.26) with the choice (1.30), where $\kappa=\frac{H^{\prime}(\delta)}{H(\delta)}$ is defined by (1.29). The breather speed $c_{s}$ is given by (1.31) and $\kappa$ also defines the spatial decay rate of the breather at infinity. Since $\kappa>0$ and $\mu>0$, we have $c_{s}<c$.
@@ 4.2. Breathers in the gap $(\zeta_{2},\zeta_{1})$
For the numerical check, we plot in Figure 8 the quotient $\rho=q/p$ (real part is shown on the left panel and imaginary part is on the right panel) for the first solution (1.16) (green line) and the second solution (1.17) (magenta line). The quotient $\rho$ is computed in two different ways: by using the elliptic functions (3.15) with $\mu$ given by (3.16) and by using the explicit solutions (1.16) and (1.17). The difference is again found within the machine precision error.
Figure 8. ${\rm Re}(\rho)$ versus $x$ (left) and ${\rm Im}(\rho)$ versus $x$ (right), where $\rho=q/p$ is the quotient of the two components of eigenfunctions $\psi=(p,q)^{T}$ of the spectral problem (3.7) with the elliptic potential $\phi$ for $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$. The spectral parameter $\zeta$ is defined by (3.2) with $\gamma=\frac{1}{2}K$. The green line shows $\rho$ for the first solution (1.16) and the magenta line shows $\rho$ for he second solution (1.17).
If we proceed with the breather solution based on the Darboux transformation (1.25) with eigenfunctions (1.16) and (1.17), we get singular solutions for real $x\in\mathbb{R}$. However, based on [26, 37], we change $x\to x+iK^{\prime}$ in order to get bounded breather solutions from the singular solutions given by (1.25). In application to the elliptic wave with the profile $\phi$, the transformation $x\to x+iK^{\prime}$ leaves $\phi$ real-valued but transforms the bounded solution (1.8) into the singular solution,
@@
The breather solutions are obtained from (1.26) with the choice (1.30), where $\kappa=Z(\gamma)$ is defined by (1.29). Although $\phi$ given by (4.3) is singular for real $x\in\mathbb{R}$, the solution $u$ in (1.26) with the choice (1.30) is bounded for real $x\in\mathbb{R}$. The breather speed $c_{s}$ is given by (1.31) and $\kappa$ also defines the spatial decay rate of the breather at infinity. Since $\kappa>0$ and $\mu<0$, we have $c_{s}>c$.
@@ 5. Factorization of elliptic eigenfunctions
We have obtained eigenfunctions of the spectral problem (3.7) in the explicit form (1.16) and (1.17). The spectral parameter $\zeta\in\mathbb{R}$ is represented by the shift parameter $\gamma\in[0,K]\times[0,iK^{\prime}]$ in Lemma 1, which also uniquely parameterizes the eigenfunctions (1.16) and (1.17). This explicit representation of eigenfunctions is now compared with the factorization formulas discussed in [3] and used in other works [30, 31]. We obtain the factorization formulas in Section 5.1. Furthermore, we show that the poles and zeros of the elliptic eigenfunctions are defined by using the branch pole singularities, which cannot be unfolded with the use of elliptic functions. This is done for the simpler cases of rational and hyperbolic degenerations of the elliptic functions in Sections 5.2 and 5.3 respectively.
5.1. Derivation of the factorization formulas
By using the quotient (3.15) with a unique choice for $\mu$ in (3.16), we introduce zeros of numerators and denominators in (3.15) for $b=4\zeta_{1}\zeta_{2}\zeta_{3}$ and $c=2(\zeta_{1}^{2}+\zeta_{2}^{2}+\zeta_{3}^{2})$ in the general case with $0<\zeta_{3}<\zeta_{2}<\zeta_{1}$. In particular, we prove in Lemma 3 below that there exist exactly two symmetric pairs of four roots in $x$ labeled as $\{\pm x_{1},\pm x_{2}\}\in[-\omega,\omega]\times[-\omega^{\prime},\omega^{\prime}]$ of
$2i\zeta\left[\phi^{2}(x)+2\zeta^{2}-\zeta_{1}^{2}-\zeta_{2}^{2}-\zeta_{3}^{2}\right]+\mu=0$ (5.1)
and exactly two symmetric pairs of four roots in $x$ labeled as $\{\pm x_{1}^{*},\pm x_{2}^{*}\}\in[-\omega,\omega]\times[-\omega^{\prime},\omega^{\prime}]$ of
@@ 5.1. Derivation of the factorization formulas
where $\omega$ and $\omega^{\prime}$ are given by (2.3). Furthermore, we prove in Lemma 4 that two roots from $\{\pm x_{1},\pm x_{2}\}$ and two roots from $\{\pm x_{1}^{*},\pm x_{2}^{*}\}$ are four roots in $x$ of
$\zeta^{2}\phi(x)-\frac{i}{2}\zeta\phi^{\prime}(x)-\zeta_{1}\zeta_{2}\zeta_{3}=0$ (5.3)
in $\omega,\omega]\times[-\omega^{\prime},\omega^{\prime}]$, whereas the complement from the two sets of roots give four roots in $x$ of
@@
Poles of both numerator and denominator in (3.15) coincide with the poles of $\phi^{2}(x)$ and $\phi^{\prime}(x)$, which are double poles at the same locations $\pm\frac{v}{2}$, see (2.6) and (2.7). Since both the numerator and the denominator in each quotient of (3.15) are elliptic functions, have only four zeros in $\omega,\omega]\times[-\omega^{\prime},\omega^{\prime}]$, and their double poles coincide at $\pm\frac{v}{2}$, the factorization formula for elliptic functions (Proposition 3.3 in [3]) implies for both quotients in (3.15) that
@@
which is the quotient $\rho=\frac{q_{1}}{p_{1}}$ of the first solution $\psi=(p_{1},q_{1})^{T}$. Note that the two poles and two zeros of the elliptic function $\rho$ in (5.8) are related by
@@
We change $\mu$ to $\mu$ in (3.15) and use the same notations for the sets of roots in Lemmas 3 and 4. This yields
@@
which is the quotient $\rho=\frac{q_{2}}{p_{2}}$ of the second solution $\psi=(p_{2},q_{2})^{T}$. Note that the limit (5.12) is different from the limit (5.9) and this yields to the asymmetry between (5.10) and (5.13) in the sense that the two expressions are not obtained by replacing $\{x_{1},x_{2}\}$ with $\{x_{1}^{*},x_{2}^{*}\}$ and vice versa.
@@
$\upsilon^{2}:=\zeta_{1}^{2}+\zeta_{2}^{2}+\zeta_{3}^{2}-2\zeta^{2}-\frac{\mu}{2i\zeta}.$
In order to compute integrals explicitly, we recall the linear fractional tranformation between the elliptic profile $\phi$ and the Weierstrass’ function $\wp$ (Lemma 3.4 in [3]):
@@
Since $\phi^{2}(x)-\upsilon^{2}$ is an elliptic function with roots given by roots of (5.1), we use Lemma 3 and write it in the factorization form
@@
$s$ $:=-i\zeta+\frac{2\gamma_{1}(\upsilon^{2}\gamma_{1}\zeta^{2}-\alpha_{1}\zeta_{1}\zeta_{2}\zeta_{3})}{i\zeta(\alpha_{1}^{2}-\gamma_{1}^{2}\upsilon^{2})}+\frac{\nu H^{\prime}(\nu x_{1})}{H(\nu x_{1})}+\frac{\nu H^{\prime}(\nu x_{2})}{H(\nu x_{2})}$
$=i\frac{4\zeta^{4}-4\zeta^{2}(\zeta_{2}+\zeta_{3}-\zeta_{1})^{2}-i\zeta\mu+4\zeta_{1}\zeta_{2}\zeta_{3}(\zeta_{2}+\zeta_{3}-\zeta_{1})}{4\zeta^{3}+4\zeta(\zeta_{2}\zeta_{3}-\zeta_{1}\zeta_{2}-\zeta_{1}\zeta_{3})-i\mu}+\frac{\nu H^{\prime}(\nu x_{1})}{H(\nu x_{1})}+\frac{\nu H^{\prime}(\nu x_{2})}{H(\nu x_{2})}.$ (5.19)
By using the quotient (5.10) for $\rho$, we hence obtain the explicit formula for $q_{1}=p_{1}\rho$ as
@@
we obtain the explicit expressions for the first solution in the form:
@@
We replace $\mu$ by $\mu$ in (5.15). This replaces $\upsilon^{2}$ with
@@
Since $\phi^{2}(x)-(\upsilon^{*})^{2}$ is an elliptic function with roots given by roots of (5.2), we use Lemma 3 and write it in the factorization form
@@
we obtain the explicit expressions for the second solution in the form:
$\left(\begin{matrix}p_{2}\\ q_{2}\end{matrix}\right)=\left(\begin{matrix}\frac{H(\nu(x-x_{1}^{*}))H(\nu(x-x_{2}^{*}))}{\Theta(\nu x-\alpha)\Theta(\nu x+\alpha)\Theta(\alpha+\nu x_{1}^{*})\Theta(\alpha+\nu x_{2}^{*})}e^{\frac{i\pi\nu}{2K}(x_{1}^{*}+x_{2}^{*})}\\ \frac{H(\nu(x+x_{1}))H(\nu(x+x_{2}))}{\Theta(\nu x-\alpha)\Theta(\nu x+\alpha)\Theta(\alpha-\nu x_{1})\Theta(\alpha-\nu x_{2})}e^{-\frac{i\pi\nu}{2K}(x_{1}+x_{2})}\end{matrix}\right)e^{-sx}.$ (5.26)
@@
Thi relation follows by taking the limit $\rho(x)$ in (5.10) as $x\to-\nu^{-1}(iK^{\prime}+\alpha)$ or $\rho(x)$ in (5.11) as $x\to\nu^{-1}(iK^{\prime}+\alpha)$:
@@
where we have used Lemma 4.2 in [3] for the second equality. This agrees with (4.1) and (4.2) for $\delta=2\alpha$. By using (1.19), we substitute values of $x_{1}$, $x_{2}$, $x_{1}^{*}$, $x_{2}^{*}$ into (5.21) and obtain

### v2

@@ 3.2. Characteristic polynomial for traveling waves
where all functions depend on only one variable which stands for the traveling wave coordinate $x+ct$.
The following lemma gives the admissible values of $\mu$, where we use Lemma 4 and a relation to Weierstrass’ function $\wp$.
Lemma 5.
Let $a\in[0,\omega]\times[0,\omega^{\prime}]$ be defined by (3.5) and (3.6). Then, $\mu^{2}=4(\wp^{\prime}(a))^{2}$.
Proof.
Since (3.12) is the linear algebraic system, the admissible values of $\mu$ are found from the characteristic equation
$\left|\begin{array}[]{ll}4i\zeta^{3}+2i\zeta\phi^{2}-ic\zeta-\mu&4\zeta^{2}\phi-2i\zeta\phi^{\prime}+2\phi^{3}-\phi^{\prime\prime}-c\phi\\ 4\zeta^{2}\phi+2i\zeta\phi^{\prime}+2\phi^{3}-\phi^{\prime\prime}-c\phi&-4i\zeta^{3}-2i\zeta\phi^{2}+ic\zeta-\mu\end{array}\right|=0,$ (3.13)
whereas the quotient $\rho=q/p$ for the eigenfunction $\psi=(p,q)^{T}$ satisfies
@@
due to the parameterization (1.7). Substituting (2.15) and (3.6) into (3.15) and using (2.2), we obtain
@@
where we used the first-order quadrature $(\wp^{\prime})^{2}=4(\wp-e_{1})(\wp-e_{2})(\wp-e_{3})$ for Weierstrass’ elliptic function $\wp$. ∎
@@ 3.3. Explicit expression for $\mu$
By Lemma 5, there exist two admissible roots for $\mu$ for the two linearly independent solutions of the spectral problem (3.7). The following lemma shows that $\mu=-2\wp^{\prime}(a)$ is uniquely selected for the two solutions (1.16) and (1.17) in Theorem 1.
Lemma 6.
For the eigenfunctions of the Lax system (1.2) defined by (1.15) (1.16) and (1.17), the value of $\mu$ is given by
$\mu=-4\nu^{3}k^{2}\operatorname{sn}(\gamma)\operatorname{cn}(\gamma)\operatorname{dn}(\gamma),$ (3.16)
Proof.
We note from (1.16) that $p_{1}(x_{0})=q_{1}(x_{0})$ at
@@
where we have used expressions (2.4), (2.5), (2.6), (2.7), and (3.6). By using [17, 8.177],
@@
we obtain $\mu=-2\wp^{\prime}(a)$. Since $a=(\gamma+iK^{\prime})/\nu$, we use (1.22) and (2.1) to rewrite the expression for $\mu$ in the form:
$\mu=-2\wp^{\prime}(a)=\frac{4\nu^{3}\operatorname{cn}(\nu a)\operatorname{dn}(\nu a)}{\operatorname{sn}^{3}(\nu a)}=-4\nu^{3}k^{2}\operatorname{sn}(\gamma)\operatorname{cn}(\gamma)\operatorname{dn}(\gamma),$
which yields (3.16). ∎
The explicit expression (3.16) enables a unique parameterization of the square root of the characteristic polynomial $P(\zeta)$ in the definition (3.15) of $\mu$. We give explicit expressions for $\mu$ in the gaps $(\zeta_{2},\zeta_{1})$ and $(0,\zeta_{3})$ to show that the different branches of the square root of $P(\zeta)$ are chosen in between the two gaps.
Example 3.
@@
due to (2.11). For $\zeta\in(0,\zeta_{3})$, we write $\gamma=\delta+iK^{\prime}$ with $\delta\in(2\alpha,K)$. The expression (3.16) can be rewritten as
@@
since the dispersion relation (2.11) can be rewritten for $\gamma=\delta+iK^{\prime}$ as
@@ 4. Breathers on the elliptic wave background
We use a linear combination of the two eigenfunctions in Theorem 1 in order to construct breathers by using the Darboux transformation (1.25). This leads to the expression (1.26). We intend here to show for the gaps $(0,\zeta_{3})$ and $(\zeta_{2},\zeta_{1})$ that the choice (1.30) produces a bounded and real-valued solution $u=u(x,t)$ of the mKdV equation (1.1). Sections 4.1 and 4.2 report the corresponding results in the two gaps, from which breather solutions in Figures 34 and 56 are constructed.
@@ 4.1. Breathers in the half gap $(0,\zeta_{3})$
Figure 7. ${\rm Re}(\rho)$ versus $x$ (left) and ${\rm Im}(\rho)$ versus $x$ (right), where $\rho=q/p$ is the quotient of the two components of eigenfunctions $\psi=(p,q)^{T}$ of the spectral problem (3.7) with the elliptic potential $\phi$ for $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$. The spectral parameter $\zeta$ is defined by (3.4) with $\delta=\frac{1}{2}(2\alpha+K)$. The green line shows $\rho$ for the first solution (4.1) and the magenta line shows $\rho$ for the second solution (4.2).
For the numerical check, we plot in Figure 7 the quotient $\rho=q/p$ (real part is shown on the left panel and imaginary part is on the right panel) for the first solution (4.1) (green line) and the second solution (4.2) (magenta line). The quotient $\rho$ is computed in two different ways: by using the elliptic functions (3.14) with $\mu$ given by (3.18) and by using the explicit solutions (4.1) and (4.2). The two representations are mathematically equivalent and hence the computational difference between them is found within the machine precision error.
The breather solutions are obtained from (1.26) with the choice (1.30), where $\kappa=\frac{H^{\prime}(\delta)}{H(\delta)}$ is defined by (1.29). The breather speed $c_{s}$ is given by (1.31) and the spatial decay rate of the breather at infinity is defined by $\kappa$. Since $\kappa>0$ and $\mu>0$, we have $c_{s}<c$.
@@ 4.2. Breathers in the gap $(\zeta_{2},\zeta_{1})$
For the numerical check, we plot in Figure 8 the quotient $\rho=q/p$ (real part is shown on the left panel and imaginary part is on the right panel) for the first solution (1.16) (green line) and the second solution (1.17) (magenta line). The quotient $\rho$ is computed in two different ways: by using the elliptic functions (3.14) with $\mu$ given by (3.16) and by using the explicit solutions (1.16) and (1.17). The computationa difference between the two equivalent representations is again found within the machine precision error.
Figure 8. ${\rm Re}(\rho)$ versus $x$ (left) and ${\rm Im}(\rho)$ versus $x$ (right), where $\rho=q/p$ is the quotient of the two components of eigenfunctions $\psi=(p,q)^{T}$ of the spectral problem (3.7) with the elliptic potential $\phi$ for $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$. The spectral parameter $\zeta$ is defined by (3.2) with $\gamma=\frac{1}{2}K$. The green line shows $\rho$ for the first solution (1.16) and the magenta line shows $\rho$ for the second solution (1.17).
If we proceed with the breather solution based on the transformation formula (1.26) with eigenfunctions (1.16) and (1.17), we get singular solutions for real $x\in\mathbb{R}$. However, based on [29, 41], we change $x\to x+iK^{\prime}$ in order to get bounded breather solutions. In application to the elliptic wave with the profile $\phi$, the transformation $x\to x+iK^{\prime}$ leaves $\phi$ real-valued but transforms the bounded solution (1.8) into the singular solution,
@@
The breather solutions are obtained from (1.26) with the choice (1.30), where $\kappa=Z(\gamma)$ is defined by (1.29). Although $\phi$ given by (4.3) is singular for real $x\in\mathbb{R}$, the solution $u$ in (1.26) with the choice (1.30) is bounded for real $x\in\mathbb{R}$. The breather speed $c_{s}$ is given by (1.31) and the spatial decay rate of the breather at infinity is defined by $\kappa$. Since $\kappa>0$ and $\mu<0$, we have $c_{s}>c$.
@@ 5. Factorization of elliptic eigenfunctions
We have obtained eigenfunctions of the spectral problem (3.7) in the explicit form (1.16) and (1.17). The spectral parameter $\zeta\in\mathbb{R}$ is represented by the shift parameter $\gamma\in[0,K]\times[0,iK^{\prime}]$ from the dispersion relation (1.14), which also parameterizes the eigenfunctions (1.16) and (1.17). This explicit representation of eigenfunctions is now compared with the factorized form discussed in [3] and used in other works [33, 34]. We obtain the factorized form in Section 5.1. Furthermore, we show that the poles and zeros of the elliptic eigenfunctions depend on the spectral parameter via the branch point singularities, which cannot be unfolded by using elliptic functions. This is done for the simpler cases of rational and hyperbolic degenerations of the elliptic functions in Sections 5.2 and 5.3 respectively.
5.1. Derivation of the factorized forms
By using the quotient (3.14) with $b=4\zeta_{1}\zeta_{2}\zeta_{3}$ and $c=2(\zeta_{1}^{2}+\zeta_{2}^{2}+\zeta_{3}^{2})$, and a unique choice for $\mu$ in (3.16), we introduce zeros of numerators and denominators in the general case $0<\zeta_{3}<\zeta_{2}<\zeta_{1}$. In particular, we prove in Lemma 7 below that there exist exactly two symmetric pairs of four roots in $x$ labeled as $\{\pm x_{1},\pm x_{2}\}$ in $\omega,\omega]\times[-\omega^{\prime},\omega^{\prime}]$ for the numerator of the first quotient in (3.14),
$2i\zeta\left[\phi^{2}(x)+2\zeta^{2}-\zeta_{1}^{2}-\zeta_{2}^{2}-\zeta_{3}^{2}\right]+\mu=0,$ (5.1)
and exactly two symmetric pairs of four roots in $x$ labeled as $\{\pm x_{1}^{*},\pm x_{2}^{*}\}$ in $\omega,\omega]\times[-\omega^{\prime},\omega^{\prime}]$ for the denominator of the second quotient in (3.14),
@@ 5.1. Derivation of the factorized forms
where $\omega$ and $\omega^{\prime}$ are given by (2.3). Furthermore, we prove in Lemma 8 that two roots from $\{\pm x_{1},\pm x_{2}\}$ and two roots from $\{\pm x_{1}^{*},\pm x_{2}^{*}\}$ are four roots in $x$ for the denominator of the first quotient in (3.14),
$\zeta^{2}\phi(x)-\frac{i}{2}\zeta\phi^{\prime}(x)-\zeta_{1}\zeta_{2}\zeta_{3}=0,$ (5.3)
in $\omega,\omega]\times[-\omega^{\prime},\omega^{\prime}]$, whereas the complement from the two sets of roots give four roots in $x$ for the numerator of the second quotient in (3.14),
@@
With the use of Lemmas 7 and 8, we obtain the factorized form for the eigenfunctions of the spectral problem (1.3). The factorized form provides an alternative definition of the eigenfunctions compared to the representations (1.16) and (1.17) in Theorem 1. The corresponding result is given by the following theorem.
Theorem 2.
Let $u(x,t)=\phi(x+ct)$ be defined by (1.8) for $0<\zeta_{3}<\zeta_{2}<\zeta_{1}$. Define $\{x_{1},x_{2}\}\in[-\omega,\omega]\times[-\omega^{\prime},\omega^{\prime}]$ and $\{x_{1}^{*},x_{2}^{*}\}\in[-\omega,\omega]\times[-\omega^{\prime},\omega^{\prime}]$ by roots of (5.1) and (5.2) respectively, which are simultaneously roots of (5.3). There exist two linearly independent solutions of the Lax system (1.2) in the form (1.15) with
and
$\left(\begin{matrix}p_{2}(x)\\ q_{2}(x)\end{matrix}\right)=\left(\begin{matrix}\frac{H(\nu(x-x_{1}^{*}))H(\nu(x-x_{2}^{*}))}{\Theta(\nu x-\alpha)\Theta(\nu x+\alpha)\Theta(\alpha+\nu x_{1}^{*})\Theta(\alpha+\nu x_{2}^{*})}e^{\frac{i\pi\nu}{2K}(x_{1}^{*}+x_{2}^{*})}\\ \frac{H(\nu(x+x_{1}))H(\nu(x+x_{2}))}{\Theta(\nu x-\alpha)\Theta(\nu x+\alpha)\Theta(\alpha-\nu x_{1})\Theta(\alpha-\nu x_{2})}e^{-\frac{i\pi\nu}{2K}(x_{1}+x_{2})}\end{matrix}\right)e^{-sx},$ (5.9)
where
$s=i\frac{4\zeta^{4}-4\zeta^{2}(\zeta_{2}+\zeta_{3}-\zeta_{1})^{2}-i\zeta\mu+4\zeta_{1}\zeta_{2}\zeta_{3}(\zeta_{2}+\zeta_{3}-\zeta_{1})}{4\zeta^{3}+4\zeta(\zeta_{2}\zeta_{3}-\zeta_{1}\zeta_{2}-\zeta_{1}\zeta_{3})-i\mu}+\frac{\nu H^{\prime}(\nu x_{1})}{H(\nu x_{1})}+\frac{\nu H^{\prime}(\nu x_{2})}{H(\nu x_{2})}.$ (5.10)
Proof.
The proof is computational and consists of the following steps.
@@
Poles of both numerator and denominator in (3.14) coincide with the poles of $\phi^{2}(x)$ and $\phi^{\prime}(x)$, which are double poles at the same locations $\pm\frac{v}{2}$, see (2.6) and (2.7). Since both the numerator and the denominator in each quotient of (3.14) are elliptic functions, which have four zeros in $\omega,\omega]\times[-\omega^{\prime},\omega^{\prime}]$ and two double poles at $\pm\frac{v}{2}$, the factorized form for elliptic functions, see Proposition 3.3 in [3], implies for both quotients in (3.14) that
@@
which is the quotient $\rho=\frac{q_{1}}{p_{1}}$ of the first solution $\psi=(p_{1},q_{1})^{T}$ of the spectral problem (3.7). The two poles and two zeros of the elliptic function $\rho$ in (5.11) are related by
@@
We change $\mu$ to $\mu$ in (3.14) and use the same notations for the sets of roots in Lemmas 7 and 8. This yields
@@
which is the quotient $\rho=\frac{q_{2}}{p_{2}}$ of the second solution $\psi=(p_{2},q_{2})^{T}$ of the spectral problem (3.7). The limit (5.15) is different from the limit (5.12) and this yields to the asymmetry between (5.13) and (5.16) in the sense that the two expressions are not obtained by replacing $\{x_{1},x_{2}\}$ with $\{x_{1}^{*},x_{2}^{*}\}$ and vice versa.
@@
$\upsilon^{2}:=\zeta_{1}^{2}+\zeta_{2}^{2}+\zeta_{3}^{2}-2\zeta^{2}-\frac{\mu}{2i\zeta}.$ (5.19)
In order to compute integrals explicitly, we recall the linear fractional tranformation between the elliptic profile $\phi$ and the Weierstrass’ function $\wp$, see Lemma 3.4 in [3]:
@@
Since $\phi^{2}(x)-\upsilon^{2}$ is an elliptic function with roots given by roots of (5.1), we use Lemma 7 and write it in the factorized form
@@
$s:=-i\zeta+\frac{2\gamma_{1}(\upsilon^{2}\gamma_{1}\zeta^{2}-\alpha_{1}\zeta_{1}\zeta_{2}\zeta_{3})}{i\zeta(\alpha_{1}^{2}-\gamma_{1}^{2}\upsilon^{2})}+\frac{\nu H^{\prime}(\nu x_{1})}{H(\nu x_{1})}+\frac{\nu H^{\prime}(\nu x_{2})}{H(\nu x_{2})},$
which yields (5.10) from (5.19) and (5.20). By using the quotient (5.13) for $\rho$, we hence obtain the explicit formula for $q_{1}=p_{1}\rho$ as
@@
we obtain the explicit expressions for the first solution in the form (5.8).
@@
We replace $\mu$ by $\mu$ in (5.18). This replaces $\upsilon^{2}$ in (5.19) with
@@
Since $\phi^{2}(x)-(\upsilon^{*})^{2}$ is an elliptic function with roots given by roots of (5.2), we use Lemma 7 and write it in the factorized form
@@
we obtain the explicit expressions for the second solution in the form (5.9). ∎
It remains to prove that $s^{*}=-s$, which we do in the following lemma.
@@
This relation follows by taking the limit $\rho(x)$ in (5.13) as $x\to-\nu^{-1}(iK^{\prime}+\alpha)$ or $\rho(x)$ in (5.14) as $x\to\nu^{-1}(iK^{\prime}+\alpha)$:
@@
where we have used Lemma 4.2 in [3] for the second equality. This agrees with (4.1) and (4.2) for $\delta=2\alpha$. We substitute values of $x_{1}$, $x_{2}$, $x_{1}^{*}$, $x_{2}^{*}$ into (5.8) and obtain